Monthly food expenditures for families in Caloocan City average Php 4 500 with a standard deviation of Php 750. Assuming that the monthly food expenditures are normally distributed, what percentages of the expenditures are between Php 3 900 and Php 5 850?


1
Expert's answer
2022-02-15T17:13:38-0500

"P(3900<X<5850)=P(\frac{3900-4500}{750}<Z<\frac{5850-4500}{750})="

"P(-0.8<z<1.8)=P(Z<1.8)-P(Z<-0.8)=0.5722=75.22\%."